OpenStudy (anonymous):

Prove the Trigonometric Equation sinx/1-cosx +sinx/1+cosx =2cscx

OpenStudy (anonymous):

I currently have sinx + sinx/1+cosx

OpenStudy (anonymous):

\[\Large \frac{ \sin \theta }{ 1-\cos \theta }+\frac{\sin \theta}{1+\cos \theta}=2\csc \theta\] correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Alright, does it matter if you make the RHS look like the LHS or do you have to simplify the LHS to the RHS

OpenStudy (anonymous):

I have to simplify the RHS to the LHS because it's more complex

OpenStudy (anonymous):

I mean LHS to the RHS

OpenStudy (anonymous):

Alright, well a tip is to work on both sides, get one side simplified to a point and then have the other side "catch up" to it. Anyways this looks simple enough, we just have to go and add the fractions so by getting like denominators we get \[\Large \frac{\sin \theta(1+\cos \theta)+\sin \theta(1-\cos \theta)}{(1-\cos \theta)(1+\cos \theta)}\]

OpenStudy (anonymous):

I'm lost

OpenStudy (anonymous):

Alright, well you know to add \[\Large \frac{a}{b}+\frac{c}{d}\] that you have to get both fractions with like denominators correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then you can add the numerators

OpenStudy (anonymous):

Never mind I got you

OpenStudy (anonymous):

OK, well committ this to memory cause it makes these^^^ a lot faster \[\Large \frac{a}{b}+\frac{c}{d}=\frac{a \cdot d}{b \cdot d}+\frac{c \cdot b}{d \cdot b}=\frac{ad+cb}{db}\]

OpenStudy (anonymous):

Okay sorry, it's just looking at that makes me confused carry on

OpenStudy (anonymous):

Yeah just take the numerator of the first and multiply it by the denomin of the second then do that for the other fraction and then multiply the denominators, it saves time in simplifying them and then adding them. Anyways we get \[ \Large \frac{\sin \theta+\sin \theta \cos \theta+\sin \theta-\sin \theta\cos \theta}{1-\cos^2 \theta}\]

OpenStudy (anonymous):

\[\Large \frac{2\sin \theta}{\sin^2 \theta}\]

OpenStudy (anonymous):

Yeah, from there I got 2sin^2x/1-cos^2x

OpenStudy (anonymous):

2sin theta/sin ^2 theta

OpenStudy (anonymous):

I think you can simplify from there

OpenStudy (anonymous):

2*1/sin theta

OpenStudy (anonymous):

I'm sorry I don't know what to do next

OpenStudy (anonymous):

1/sin theta=csc theta

OpenStudy (anonymous):

The sin in the numerator cancels out with one of the sin in the denom.

OpenStudy (anonymous):

1-cos^2 theta=sin^2 theta use this identity in the denominator

OpenStudy (anonymous):

\[\Large \frac{2\sin \theta}{\sin^2 \theta}= \frac{2\sin \theta}{\sin \theta \cdot \sin \theta}=\frac{2}{\sin \theta}=2 \cdot \frac{1}{\sin \theta}= 2\csc \theta\]

OpenStudy (anonymous):

Thanks a bunch dude

OpenStudy (anonymous):

Yw